### v4

# Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training

Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. We then develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training. Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters. Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. We then develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training. Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters. Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
[…]
Figure 1: Illustration of the three primary challenges addressed in this work, with the third being the most crucial.
1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" [25]. However, the report provides little detail on the methodology behind this prediction approach.
This technique is critical for the pre-training of large language models (LLMs). On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a highly challenging endeavor [14, 39, 8]. The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste. On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage [12, 21].
In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge [29]. This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions. Meanwhile, we encounter three major challenges (Figure 1):
Challenge1: Limited access to and understanding of pre-training corpora. Currently, few open-source foundation models fully disclose their pre-training datasets. Pre-training from scratch is prohibitively expensive, requiring massive data collection and compute. For example, Meta reports that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs over 100 million dollars [8].
Challenge2: Limitations of current evaluation methods for pre-trained models. Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining [29, 36]. However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation [24, 33]. These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge. In particular, assessing a model’s knowledge using a single query template is unreliable [40].
Challenge3: Limitations of frequency-based metrics in predicting model performance. The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data [1, 37]. Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data [16, 22, 2]. However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends. For example, despite similar frequencies of occurrence, the model may struggle to pinpoint the location of Apple’s headquarters due to the pervasive nature of the subject. In contrast, identifying NVIDIA’s headquarters might be relatively straightforward. Furthermore, frequency-based methods do not consider the impact of model size.
Our work primarily focuses on effectively addressing these three challenges to predict model knowledge from pre-training signals, thereby estimating the model’s performance on downstream tasks without any additional training. Specifically:
First, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained LLMs. With comprehensive access to the full pre-training data, we are able to systematically trace and examine the knowledge that each model was exposed to during pre-training, enabling a detailed understanding of the relationship between pre-training data and downstream performance.
Second, inspired by the evaluation methods for fine-tuning LLMs [40], we develop a robust multimodal evaluation approach. Recognizing that the same knowledge can correspond to various questioning styles, each style allows for the design of corresponding templates. For each category of question-answering tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
Finally, we propose SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training. Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention. Across various model sizes, ranging from 14 million to 176 billion parameters, SMI significantly outperforms the co-occurrence-based baseline. For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
Moreover, the comparative analysis across models shows diminishing returns from both scaling model size and optimizing data. As model size increases, its marginal impact on QA performance decreases, suggesting that larger capacity alone cannot guarantee continued gains. Likewise, while higher SMI values are generally associated with better accuracy, pushing SMI further becomes increasingly difficult due to its logarithmic growth. Notably, even in the largest models within each family, QA accuracy consistently plateaus around 80% when SMI is maximized. This suggests a fundamental limitation in how much factual knowledge can be internalized during pre-training, highlighting the need for external augmentation methods such as retrieval or prompting in downstream stages.
In summary, our contributions are threefold:
  1. 1.
    We propose an information theory-based method and introduce the SMI metric, which accurately predicts the knowledge retention of pre-trained LLMs using only pre-training signals, without requiring any additional training.
  2. 2.
    Across LLMs with varying sizes from 14 million to 176 billion parameters, the SMI metric exhibits a strong linear relationship with QA accuracy, significantly outperforms the traditional co-occurrence-based baseline.
  3. 3.
    Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns. Even under optimal conditions, QA accuracy plateaus around 80%, reflecting the intrinsic limit of knowledge that can be memorized during pre-training.
2 Related Work
Pre-training data and knowledge retention
The characteristics of pre-training data are crucial for knowledge retention and downstream performance in LLMs. Many studies emphasize the role of repetition frequency. For example, Carlini et al. [6] identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al. [7] showed that sequences repeated over 500 times are completed with over 40% accuracy. Ju et al. [15] found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning. Allen-Zhu and Li [2] suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter. Additionally, Razeghi et al. [28] and Yadlowsky et al. [38] linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al. [23] associated task prevalence in pre-training with better performance on tasks like ROT13.
Model size and knowledge retention
Furthermore, the model size significantly influences the knowledge retention of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern [17].  Kandpal et al. [16] evaluate the BLOOM model families [30] on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size [19].
Knowledge triples as an evaluation tool
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. [27] employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. [13], Ju et al. [15] further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. [37] and Allen-Zhu and Li [1] use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
3 Methodology
To address the three challenges outlined in Section 1, we propose a method to predict model knowledge solely from pre-training signals. We begin with prior work and explore the use of the co-occurrence metric. Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes. Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
3.1 Starting from the Co-occurrence Metric
Co-occurrence metric baseline
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns [16, 22, 2]. Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach [16]. There is a linear relationship between QA accuracy and the logarithm of the number of documents [16]. The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
Knowledge triples
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object [37]. For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
$F(q_{s,r})=o,\quad(q\in\mathbf{Q},o\in\mathbf{O}).$ (1)
Here, $q_{s,r}$ represents the combined representation of the subject $s$ and the relation $r$, while $o$ denotes the representation of the object. For instance, if $s=\text{Apple}$ and $r=\text{headquarter}$, then $q_{s,r}$ can be expressed as statements like “The headquarters of Apple is in…”, “Apple is headquartered in…”, or “Apple’s head office is based in…”. Similarly, $o$ corresponds to object representation, such as “Cupertino”, “Cupertino, California”, or “Cupertino city”. The sum of all $q_{s,r}$ representations is denoted as $\mathbf{Q}$, and the sum of all $o$ representations is denoted as $\mathbf{O}$.
Multi-template evaluation method
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs [40], we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
Specifically, by constructing synonymous sentences, we generate a substantial number of query templates that are semantically similar but vary in form. For each type of knowledge triple, we produce 20 templates that differ in length and structure. Each template $q$ represents a specific instance within the query set $\mathbf{Q}$, and these 20 selected $q$ templates approximate the entirety of $\mathbf{Q}$.
Linear relationship
On the one hand, we calculate metrics from the pre-training data. Each QA task consists of a subject and an object. We perform document-level retrieval on the pre-training data and compute three values: $P(s)$ represents the proportion of documents containing the subject out of the total number of documents $\mathbf{N}$ in the corpus, $P(o)$ denotes the proportion of documents containing the object out of $\mathbf{N}$, and $P(s,o)$ signifies the proportion of documents containing both the subject and the object out of $\mathbf{N}$. The co-occurrence metric is given by $\log_{2}(P(s,o))$.
On the other hand, by employing the multi-template evaluation approach, we measure the model’s accuracy on QA tasks. To examine the relationship between the co-occurrence metric and QA accuracy, we fit a linear regression model and compute the R² along with the Mean Squared Error (MSE). These metrics allow us to evaluate both the strength and the stability of the linear relationship.
3.2 Considering the Knowledge Specificity to Obtain the MI Metric
Knowledge specificity
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities [31]. Therefore, we aim to use MI to improve the co-occurrence metric.
MI metric
[…]
$I(s,o)$ $=P(s,o)\log\left(\frac{P(s,o)}{P(s)P(o)}\right),$ (2)
$MI(s,o)$ $=\log_{2}(I(s,o)).$ (3)
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$ (Figure 2), we take its logarithm to obtain the MI metric.
Figure 2: Evaluation data statistics. Left: Co-occurrence frequencies of QA entity pairs in The Pile corpus. Right: Distribution of 12,671 QA tasks across 15 relation types.
3.3 Considering the Model Size to Obtain the Size-dependent MI metric
Model size
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger [17]. In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters [16]. Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
The method of integrating model size is not unique. A guiding principle is that, for the same data, a higher $\Phi$ should correspond to a higher value of the new metric, reflecting improved QA performance. However, the gain in accuracy should diminish progressively. Importantly, as $\Phi$ approaches infinity, the new metric should converge to the MI metric, indicating that the model’s memory capacity is sufficient to fully capture the relevant knowledge.
Size-dependent MI metric
Based on these principles, we propose the SMI metric—an exponential function with the normalized MI as the base and $1+\frac{1}{\Phi}$ as the exponent.
$SMI(s,o,\Phi)=Norm(MI(s,o))^{1+\frac{1}{\Phi}}.$ (4)
In the above formula, the MI metric quantifies the information of knowledge triples in the pre-training data, while the model size $\Phi$ implies the memory capacity of LLMs. These two factors jointly determine the model’s knowledge retention. We normalize the MI metric from its original range of -30 to 0 onto a 0–1 scale. Since the normalized MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
Figure 3: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families. Model size is reported in billions of parameters and plotted on a base-2 logarithmic scale. Variations in pre-training data across model families may account for differences in performance.
Table 1: R² and MSE (×$\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters. Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
Model R² ($\uparrow$ higher is better) MSE (×$\times 10^{-2}$, $\downarrow$ lower is better)
CO-OCCUR MI SMI CO-OCCUR MI SMI
BLOOM
bloom-560m 0.384 0.589 0.624 8.566 7.729 7.652
bloom-1b1 0.577 0.715 0.756 9.123 7.950 7.824
bloom-1b7 0.584 0.740 0.772 9.389 8.083 7.970
bloom-3b 0.705 0.782 0.807 9.314 7.863 7.764
bloom-7b1 0.676 0.804 0.814 9.943 8.309 8.263
bloom (176b) 0.644 0.764 0.764 7.327 6.029 6.029
GPT-Neo
gpt-neo-125m 0.349 0.531 0.593 7.004 7.107 6.999
gpt-neo-1.3B 0.671 0.810 0.859 8.233 7.444 7.265
gpt-neo-2.7B 0.672 0.826 0.850 8.168 7.320 7.233
gpt-j-6b 0.631 0.862 0.869 9.026 7.720 7.692
gpt-neox-20b 0.593 0.841 0.843 8.557 7.476 7.468
Pythia
pythia-14m 0.240 0.375 0.496 6.855 7.183 6.978
pythia-31m 0.287 0.381 0.549 6.905 7.202 6.905
pythia-70m 0.355 0.430 0.585 6.749 7.051 6.783
pythia-160m 0.454 0.522 0.611 6.877 6.969 6.796
pythia-410m 0.524 0.701 0.759 7.439 7.035 6.889
pythia-1.4b 0.574 0.799 0.835 8.254 7.325 7.202
pythia-2.8b 0.631 0.845 0.866 8.593 7.505 7.421
pythia-6.9b 0.637 0.855 0.865 8.538 7.373 7.334
pythia-12b 0.602 0.853 0.858 8.741 7.514 7.494
TinyLlama
TinyLlama-1.1B 0.641 0.760 0.839 7.803 7.135 6.887
Ours
ours-1.6b 0.710 0.794 0.848 8.563 7.696 7.462
ours-7b 0.784 0.847 0.855 8.983 8.033 7.997
ours-13b 0.661 0.816 0.820 8.585 7.688 7.674
4 Experiments
In the Experimental Setup section, we begin by cleaning and enhancing an open-source Wikipedia knowledge triple dataset to make it suitable for inference with pre-trained models. We then introduce five large-scale open-source pre-training corpora and perform retrieval-based analyses on them. This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol. In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across three evaluation metrics. We also conduct a hyperparameter study to further explore model behavior under different configurations.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours [9]. We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
We conduct a document-level retrieval analysis of the following pre-training datasets, selected based on their usage in the LLMs under our consideration:
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality [11].
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency [18, 30].
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs [32].
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM [26].
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency [10].
Pre-training data retrieval
We process the pre-training corpus at the document level. For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences. Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple. Subsequently, we compute the SMI metric based on the MI metric and the model size. The retrieval process runs on 100 CPUs for 3 days.
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs [34]. These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion [30].
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile [4, 5, 35].
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable [3].
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions [41, 32, 20]. The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
Evaluation setting
We assess these pre-trained models using the designated evaluation dataset. During the evaluation process, we set the temperature to 0.7, the maximum new tokens to 32, and the precision to bfloat16. We consider a response to be correct if the object appears within it. To ensure stability in the evaluation, we perform inference 20 times for each of the 20 queries in the QA task, subsequently calculating the average accuracy of 400 responses. This inference process runs for 5 days on 8 H20 GPUs.
The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2). The majority of QA data co-occur between 10 and 1000 times. To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values. For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2. The SMI metric is computed based on the MI metric.
Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship. Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
(a) bloom (176b)
(b) gpt-neox-20b
(c) pythia-12b
(d) ours-13b
Figure 4: Relationship between the SMI metric and average accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
(a) $\text{R}^{2}$ on smaller models
(b) $\text{R}^{2}$ on larger models
Figure 5: Comparison of $\text{R}^{2}$ results for three metrics—CO-OCCUR, MI, and SMI—across models of different sizes. (a) Smaller models (bloom-560m, gpt-neo-1.3B, pythia-14m, ours-1.6b); (b) Larger models (bloom, gpt-neox-20b, pythia-12b, ours-13b). $\uparrow$ Higher is better.
Table 2: Hyperparameter analysis of the SMI metric across different BLOOM models. (a) illustrates the impact of varying unit scales, while (b) shows the effect of different pre-normalization ranges.
(a) SMI values across varying unit scales
Unit bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) $10^{8}$ 0.590 0.716 0.741 0.783 0.804 0.764 $10^{9}$ 0.600 0.723 0.745 0.785 0.805 0.764 $10^{10}$ 0.624 0.756 0.772 0.807 0.814 0.764 $10^{11}$ 0.394 0.497 0.586 0.760 0.826 0.766 $10^{12}$ 0.343 0.222 0.197 0.235 0.349 0.769
(b) SMI values under different pre-normalization ranges
Range bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) [30,0][-30,0][ - 30 , 0 ] 0.624 0.756 0.772 0.807 0.814 0.764 [35,0][-35,0][ - 35 , 0 ] 0.626 0.750 0.766 0.800 0.811 0.764 [40,0][-40,0][ - 40 , 0 ] 0.625 0.746 0.761 0.797 0.809 0.764 [45,0][-45,0][ - 45 , 0 ] 0.623 0.742 0.758 0.794 0.808 0.764 [50,0][-50,0][ - 50 , 0 ] 0.620 0.739 0.756 0.793 0.808 0.764
4.2 Experimental Results
Results of SMI, MI, and co-occurrence metrics
The primary experimental results for the three metrics are presented in Table 1. Overall, we observe the following: (1) The SMI metric performs exceptionally well on the evaluation set, with both R² and MSE surpassing the baseline across all models. (2) For all models, the SMI metric achieves R² values greater than 0.49, and for models with parameters exceeding 1 billion, the R² exceeds 0.75. This highlights a robust linear relationship and effectively predicts QA accuracy. (3) The MI metric outperforms the baseline co-occurrence metric, while the SMI metric further exceeds the MI metric, demonstrating the effectiveness of our iterative improvement approach.
The comparison of experimental results from models of different sizes indicates diminishing returns when increasing model size. A linear relationship is observed between the models’ average accuracy on the full evaluation dataset and the logarithm of model size, suggesting that significant gains in QA performance require exponential increases in model size (Figure 3). Meanwhile, the advantage of the SMI metric over the MI metric is more pronounced in smaller models (Figure 5). As model size increases, this advantage becomes less evident. For the BLOOM 560m model, the R² values differ by 3.55E-02, and the MSE values differ by 7.75E-04. In contrast, for the BLOOM 176b model, the difference in R² values is only 2.42E-04, and the difference in MSE values is just 6.60E-06. This indicates that larger models may possess sufficient memory capacity.
Moreover, the benefits of data optimization exhibit diminishing returns and an upper limit. Although a higher SMI generally correlates with better performance, achieving further gains becomes increasingly difficult at advanced levels due to the logarithmic growth of SMI (Figure 4). Specifically, the upper accuracy limits for the models bloom, gpt-neox-20b, pythia-12b, and ours-13b are 0.638, 0.727, 0.75, and 0.771, respectively. This indicates that even with perfect optimization of the pre-training data for these models, the accuracy on the QA task cannot exceed an approximate upper limit of 0.8.
Results of hyperparameter influence on the SMI metric
The SMI framework involves two critical hyperparameters. The first is the unit for model parameter quantity, which is set by default to 1e9. The second is the pre-normalization range for the SMI metric, with a default setting of -30 to 0. We conduct ablation experiments on these two hyperparameters, and the results are presented in Table 2. From the experimental results, we observe the following: (1) The unit for model parameter quantity significantly affects the SMI results. Using 1e10 as the unit yields the best overall performance. However, as the model size increases, the optimal unit value also increases. Smaller models exhibit lower robustness to parameter variations, whereas larger models display enhanced robustness. (2) The pre-normalization range has a relatively minor effect on the SMI results. The range of -30 to 0 performs best overall. Although a smaller pre-normalization range is preferable, the minimum value in the experimental data approximates -30, suggesting that further reduction might result in data loss.
5 Limitations and Future Work
Although the SMI metric shows strong predictive power for model accuracy on closed-book QA, the current evaluation is limited to this task type. This focus is intentional, as closed-book QA serves as a direct probe of factual knowledge internalized during pre-training, rather than general reasoning or transfer ability. However, broader task coverage is necessary to assess the general applicability of SMI across diverse task formats. Some variation is also observed across model scales, suggesting that future formulations should account for size-invariance. In addition, while SMI exhibits empirical effectiveness, its mathematical foundations remain underexplored. Future work will aim to establish a more rigorous theoretical grounding by linking information-theoretic principles with LLM behavior.
6 Conclusion
In this paper, we explore how to predict model knowledge using only pre-training signals and propose effective solutions to address three major challenges. We introduce SMI, an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy without any additional training. SMI achieves $R^{2}>0.75$ on models with over one billion parameters. Experimental results further reveal that both scaling model size and optimizing data yield diminishing returns, with QA accuracy plateauing at around 80% under optimal pre-training conditions. This indicates an upper bound on factual knowledge retention during pre-training, highlighting the need for downstream augmentation such as retrieval or prompting.
This paper emphasizes the feasibility of predicting downstream performance based on pre-training signals. It offers a principled metric for data-model-task alignment, thereby providing a reference for optimizing data selection and performance prediction during the pre-training phase of LLMs. To facilitate further research, all evaluation data and analysis code are made publicly available.
Appendix A Licenses and Terms of Use
We use the following publicly available assets in this work:
We confirm that all licenses and terms of use are respected.
Table 3: Twenty templates and their corresponding accuracy values for the $\mathbf{BORN}$ relation on the ours-13b model. [S] represents the subject in the knowledge triple.
Template ACC
[S]’s birth took place in 0.186
[S] was brought into existence in 0.187
[S] was born into the world in 0.190
[S]’s nativity is 0.197
The place in which [S] was given birth to is 0.201
The hometown of [S] is 0.203
The birthplace of [S] is 0.212
[S] was delivered in 0.215
[S]’s cradle was in 0.219
[S]’s entry into the world happened in 0.225
[S] first saw the light in 0.228
[S]’s roots are in 0.230
[S]’s life began in 0.230
[S] was born and raised in 0.237
[S]’s origin is 0.238
[S] entered life in 0.251
[S] came into the world in 0.255
[S] was born in 0.255
[S]’s birth occurred in 0.263
[S]’s starting point was 0.282
Table 4: Several cases with significant deviations from the co-occurrence metric are shown here. $N_{s}$ is the number of paragraphs in which the subject appears, $N_{o}$ is the number of paragraphs in which the object appears, and $N_{s,o}$ is the number of paragraphs in which they co-occur. The first two cases exhibit rare co-occurrence but demonstrate high accuracy, whereas the latter two cases show frequent co-occurrence but exhibit low accuracy.
Subject Question Object $N_{s}$ $N_{o}$ $N_{s,o}$ $ACC$
Rare Co-occurrence but High Accuracy
Csepel SC The headquarters of Csepel SC is in Budapest 398 748,860 172 0.655
Liis Lemsalu Liis Lemsalu is a citizen of Estonia 307 967,207 160 0.570
Frequent Co-occurrence but Low Accuracy
Paris Paris is located in Europe 20,902,571 59,408,155 3,761,610 0.043
Ohio The capital of Ohio is Columbus 11,894,303 3,718,483 1,210,957 0.193
Appendix B Templates for Knowledge Triples
[…]
Appendix C Case Study
Table 4 highlights four counterintuitive cases in our pre-training data. From these, we observe the following: (1) In the first two cases, the subject and object rarely co-occur, suggesting a seemingly weak relationship. However, whenever the subject does appear, the object almost always appears alongside it, resulting in high accuracy. (2) In the last two cases, despite frequent co-occurrence between the subject and object, their individual occurrence frequencies are significantly higher, leading to low accuracy. For example, when prompted with “Paris is located in”, the model is more likely to respond with “France” rather than “Europe”.
These examples highlight the pivotal role of knowledge specificity in the memory of LLMs. It is not simply the frequencies of knowledge occurrence that enhance retention, but rather the specificity of the knowledge that proves more influential. This insight suggests that in downstream task training, even when domain-specific data is significantly fewer than the pre-training data, minimizing repetition with the pre-training data can increase the SMI, making it easier for LLMs to retain the relevant knowledge.