### v5

# Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training

Abstract
The GPT-4 technical report suggests that downstream performance can be predicted from pre-training signals, but offers little methodological detail on how to quantify this. This work address this gap by modeling knowledge retention, the capacity of a pre-trained language model to memorize factual information from its corpus, and introduce a principled method to estimate it prior to training. We propose Size-dependent Mutual Information (SMI), an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size to forecast closed-book question answering (QA) accuracy. SMI is validated through large-scale document retrieval over the disclosed pre-training corpora of 21 public and 3 custom models, combined with a robust multi-template QA evaluation. Experiments show that SMI significantly outperforms repetition-based baselines and achieves R² > 0.7 in predicting QA accuracy for models above 1B parameters, without additional training. The analysis further reveals diminishing returns from scaling data and model size and provides evidence for an intrinsic upper bound on knowledge retention achievable by pre-training alone, motivating retrieval and other augmentation strategies. The dataset and code are available at https://github.com/yuhui1038/SMI.
Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training
Changhao Jiangthanks: Equal contribution, Ming Zhang11footnotemark: 1, Yifei Cao11footnotemark: 1, Junjie Ye11footnotemark: 1, Xiaoran Fan11footnotemark: 1, Shihan Dou, Zhiheng Xi, Jiajun Sun, Yi Dong, Yujiong Shen, Jingqi Tong, Baoyu Fan, Tao Gui, Qi Zhangthanks: Corresponding author, Xuanjing Huang Fudan University, Shanghai, China IEIT Systems Co. Ltd, Shenzhen, China chjiang25@m.fudan.edu.cn, mingzhang23@fudan.edu.cn, qz@fudan.edu.cn
Figure 1: Overview of our method. The framework consists of three stages: (1) Corpus Analysis and Retrieval, where subject and object (question and answer) entities are extracted from pre-training data and their frequencies and co-occurrences are computed; (2) Information-Theoretic Metric Computation, where knowledge frequency, knowledge specificity, and model size are integrated into the size-dependent mutual information (SMI) metric; and (3) Evaluation via Closed-Book QA, where pre-trained language models are assessed using multi-template evaluation, and SMI is compared with empirical QA accuracy to analyze linear relationships and scaling effects.
1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" (OpenAI, 2023). Understanding and predicting the generalization ability of large language models (LLMs) has become a key research question in the study of scaling laws (Kaplan et al., 2020). Although prior work has established empirical relationships between loss, data size, and model parameters, these results primarily offer coarse-grained predictions of overall performance and do not address what specific factual knowledge is retained from pre-training corpora. Recent studies have shown that LLMs can memorize and regurgitate training data (Carlini et al., 2021; Biderman et al., 2023a), but these findings are largely descriptive and do not provide a systematic way to predict retention for new model–corpus pairs. Developing principled approaches for this purpose is crucial both for improving our theoretical understanding of model training dynamics and for guiding practical decisions on data curation and compute allocation (Hoffmann et al., 2022; Alabdulmohsin et al., 2022; Ruan et al., ).
This work addresses this gap by modeling knowledge retention, the capacity of a pre-trained language model to memorize factual information from its corpus, and introduces a principled method to reliably estimate it prior to training. Rather than focusing only on aggregate loss curves or perplexity, we take a more fine-grained and systematic view: we ask, given a particular fact in the pre-training corpus, what is the probability that a model will be able to accurately reproduce it in a closed-book setting after training (Petroni et al., 2019)? We term this capacity knowledge retention, and formalize it as a quantitative property of a model–corpus pair. This framing allows us to reason about model memory at the level of individual facts and categories, and to make meaningful predictions before committing to the cost of full-scale training.
We propose Size-dependent Mutual Information (SMI), an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size to forecast closed-book question answering (QA) accuracy (Roberts et al., 2020). Since pre-trained language models are optimized only for next-token prediction rather than explicit knowledge retrieval, closed-book QA provides a natural and challenging probe for assessing factual recall. Our formulation builds on the intuition that both the frequency of a fact in the corpus and its rarity influence its likelihood of being learned, and that these effects are jointly influenced by model size. Concretely, SMI captures the mutual information between knowledge subjects and objects (or questions and answers), combining these statistics with model size. The resulting predictor is simple to compute from pre-training data and scales gracefully with corpus size, making it suitable for practical use in large-scale settings.
SMI is validated through large-scale document retrieval over the disclosed pre-training corpora of 21 public and 3 custom models, combined with a robust multi-template QA evaluation. Our experimental setup reconstructs the exposure of facts in the training data, then evaluates whether models can correctly answer corresponding questions in a closed-book setting. We design multi-template QA prompts to reduce prompt variance and report results across a diverse set of model families, covering a range from 14M to 176B parameters.
Experiments show that SMI significantly outperforms repetition-based baselines and achieves the coefficient of determination (R²) > 0.7 in predicting QA accuracy for models above 1B parameters, without additional training. Relative to heuristics based solely on repetition counts, SMI yields a stronger correlation with observed QA accuracy, particularly in the large-model regime where memorization behavior appears more stable. Moreover, we observe that SMI maintains comparable predictive performance across multiple model families and domains, indicating that it captures a generalizable aspect of factual knowledge acquisition in pre-trained language models.
The analysis further reveals diminishing returns from scaling both data and model size, suggesting that pre-training alone faces an intrinsic limit in its ability to retain factual knowledge. Recent studies have similarly observed a performance ceiling on knowledge-intensive tasks, even with substantial scaling of parameters and data (Zhang et al., 2025). These observations point to inherent limitations of purely parametric approaches: beyond a certain scale, further increases in model size or training data yield only marginal gains in closed-book performance. This motivates exploring hybrid approaches that augment parametric memory with non-parametric mechanisms, including retrieval-augmented generation, external memory modules, or targeted fine-tuning, to mitigate the ceiling effects observed in pre-training. Overall, our contributions are threefold:
  1. 1.
    We formalize knowledge retention as a quantitative metric defined over model–corpus pairs, providing a new way to analyze what models actually memorize during pre-training.
  2. 2.
    We propose SMI, a data-driven and training-free predictor that integrates knowledge frequency, knowledge specificity, and model size to estimate closed-book QA accuracy.
  3. 3.
    We empirically evaluate SMI on 24 models, showing that it achieves strong predictive accuracy, characterizing diminishing returns from scaling data and model size, and suggesting the presence of an upper bound on knowledge retention achievable by pre-training alone.
2 Related Work
Pre-training data and knowledge retention
The characteristics of pre-training data are crucial for knowledge retention and downstream performance in LLMs. Many studies emphasize the role of repetition frequency. For example, Carlini et al. (2023) identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al. (2023) showed that sequences repeated over 500 times are completed with over 40% accuracy. Ju et al. (2024) found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning. Allen-Zhu and Li (2024b) suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter. Additionally, Razeghi et al. (2022) and Yadlowsky et al. (2023) linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al. (2023) associated task prevalence in pre-training with better performance on tasks like ROT13.
Model size and knowledge retention
Furthermore, the model size significantly influences the knowledge retention of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern (Kaplan et al., 2020).  Kandpal et al. (2023) evaluate the BLOOM model families (Scao et al., 2022) on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size (Li et al., 2024).
Knowledge triples as an evaluation tool
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. (2019) employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. (2024); Ju et al. (2024) further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. (2024) and Allen-Zhu and Li (2024a) use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
3 Methodology
To predict a model’s ability to recall factual knowledge before training, we develop a framework that estimates closed-book QA accuracy directly from pre-training data statistics and model size. We first revisit prior work and evaluate a simple repetition-based co-occurrence metric as a baseline. Building on information-theoretic principles, we then introduce the Mutual Information (MI) metric, which accounts for both the frequency and specificity of knowledge items and yields improved predictive performance. Finally, we extend MI to SMI, which incorporates model size as a scaling factor, providing a more accurate and scalable predictor of knowledge retention.
3.1 Starting from the Co-occurrence Metric
Co-occurrence metric baseline
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b). Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach (Kandpal et al., 2023). There is a linear relationship between QA accuracy and the logarithm of the number of documents (Carlini et al., 2023; Kandpal et al., 2023). The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
Knowledge triples
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object (Wang et al., 2024). For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
$F(q_{s,r})=o,\quad(q\in\mathbf{Q},o\in\mathbf{O}).$ (1)
Here, $q_{s,r}$ represents the combined representation of the subject $s$ and the relation $r$, while $o$ denotes the representation of the object. For instance, if $s=\text{Apple}$ and $r=\text{headquarter}$, then $q_{s,r}$ can be expressed as statements like “The headquarters of Apple is in…”, “Apple is headquartered in…”, or “Apple’s head office is based in…”. Similarly, $o$ corresponds to object representation, such as “Cupertino”, “Cupertino, California”, or “Cupertino city”. The sum of all $q_{s,r}$ representations is denoted as $\mathbf{Q}$, and the sum of all $o$ representations is denoted as $\mathbf{O}$.
Multi-template evaluation method
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs (Ye et al., 2024), we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
Specifically, by constructing synonymous sentences, we generate a substantial number of query templates that are semantically similar but vary in form. For each type of knowledge triple, we produce 20 templates that differ in length and structure. Each template $q$ represents a specific instance within the query set $\mathbf{Q}$, and these 20 selected $q$ templates approximate the entirety of $\mathbf{Q}$.
Linear relationship
On the one hand, we calculate metrics from the pre-training data. Each QA task consists of a subject and an object. We perform document-level retrieval on the pre-training data and compute three values: $P(s)$ represents the proportion of documents containing the subject out of the total number of documents $\mathbf{N}$ in the corpus, $P(o)$ denotes the proportion of documents containing the object out of $\mathbf{N}$, and $P(s,o)$ signifies the proportion of documents containing both the subject and the object out of $\mathbf{N}$. The co-occurrence metric is defined as $\log(P(s,o))$, representing the logarithm of the fraction of documents in which both the subject and object appear. For consistency with information-theoretic conventions (Shannon, 1948a), all logarithms in this paper use base 2. Since specifying the subject and object typically determines the relation, we omit the relation component for simplicity.
On the other hand, by employing the multi-template evaluation approach, we measure the model’s accuracy on QA tasks. To examine the relationship between the co-occurrence metric and QA accuracy, we fit a linear regression model and compute the R² along with the Mean Squared Error (MSE). These metrics allow us to evaluate both the strength and the stability of the linear relationship.
3.2 Considering the Knowledge Specificity to Obtain the MI Metric
Knowledge specificity
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities (Shannon, 1948b). Therefore, we aim to use MI to improve the co-occurrence metric.
MI metric
Unlike the MI in information theory, which considers all possible values of random variables, the MI metric within pre-training data considers only those instances in which the relevant entities actually appear. Given pre-training data $\mathbf{P}$ consisting of $\mathbf{N}$ paragraphs, a knowledge triple $(s,r,o)$, and corresponding $P(s)$, $P(o)$, and $P(s,o)$, the MI metric between $s$ and $o$ is defined as:
$I(s,o)$ $=P(s,o)\log\left(\frac{P(s,o)}{P(s)P(o)}\right),$ (2)
$MI(s,o)$ $=Norm(\log(I(s,o))).$ (3)
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$ (see Figure 2), we apply a logarithmic transformation followed by normalization to map its values to the range $[0,1]$, resulting in the MI metric. In rare cases where $I(s,o)$ becomes zero or negative due to data sparsity or noise, such instances are regarded as anomalies and excluded from further analysis.
Figure 2: Evaluation data statistics. Left: Co-occurrence frequencies of QA entity pairs in The Pile corpus. Right: Distribution of 12,671 QA tasks across 15 relation types.
3.3 Considering the Model Size to Obtain the Size-dependent MI metric
Model size
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger (Kaplan et al., 2020). In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters (Kandpal et al., 2023). Therefore, we propose to extend the MI metric by explicitly incorporating the parameter count $\Phi$ as a measure of model size.
The method of integrating model size is not unique. A guiding principle is that, for the same data, a higher $\Phi$ should correspond to a higher value of the new metric, reflecting improved QA performance. However, the gain in accuracy should diminish progressively. Importantly, as $\Phi$ approaches infinity, the new metric should converge to the MI metric, indicating that the model’s memory capacity is sufficient to fully capture the relevant knowledge.
Size-dependent MI metric
Since linear transformations do not alter the strength of linear relationships, an exponential formulation offers a simple yet effective way to incorporate model size while capturing non-linear effects. Based on these principles, we propose the SMI metric, an exponential function with the MI metric as the base and $1+\frac{1}{\log\Phi}$ as the exponent.
$SMI(s,o,\Phi)=MI(s,o)^{1+\frac{1}{\log\Phi}}.$ (4)
In the above formula, the MI metric quantifies the information of knowledge triples in the pre-training data, while the model size $\Phi$ implies the memory capacity of LLMs. These two factors jointly determine the model’s knowledge retention. Since the MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
4 Experiments
In the Experimental Setup section, we begin by cleaning and enhancing an open-source Wikipedia knowledge triple dataset to make it suitable for inference with pre-trained models. We then introduce five large-scale open-source pre-training corpora and perform retrieval-based analyses on them. This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol. In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across metrics.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours (Elazar et al., 2021). We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
We conduct a document-level retrieval analysis of the following pre-training datasets, selected based on their usage in the LLMs under our consideration:
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality (Gao et al., 2021).
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency (Laurençon et al., 2022; Scao et al., 2022).
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs (Soboleva et al., 2023).
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM (Penedo et al., 2023).
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency (Foundation, ).
Pre-training data retrieval
We process the pre-training corpus at the document level. For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences. Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple. Subsequently, we compute the SMI metric based on the MI metric and the model size. The retrieval process runs on 20 CPUs for 15 days.
Table 1: R² and MSE ($\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters. Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
Model R² ($\uparrow$ higher is better) MSE ($\times 10^{-2}$, $\downarrow$ lower is better)
CO-OCCUR MI SMI CO-OCCUR MI SMI
BLOOM
bloom-560m 0.384 0.589 0.591 8.566 7.729 7.723
bloom-1b1 0.577 0.715 0.719 9.123 7.950 7.939
bloom-1b7 0.584 0.740 0.743 9.389 8.083 8.072
bloom-3b 0.705 0.782 0.786 9.314 7.863 7.849
bloom-7b1 0.676 0.804 0.807 9.943 8.309 8.296
bloom 0.644 0.764 0.766 7.327 6.029 6.026
GPT-Neo
gpt-neo-125m 0.349 0.531 0.535 7.004 7.107 7.100
gpt-neo-1.3B 0.671 0.810 0.814 8.233 7.444 7.427
gpt-neo-2.7B 0.672 0.826 0.829 8.168 7.320 7.307
gpt-j-6b 0.631 0.862 0.865 9.026 7.720 7.710
gpt-neox-20b 0.593 0.841 0.843 8.557 7.476 7.469
Pythia
pythia-14m 0.240 0.375 0.378 6.855 7.183 7.177
pythia-31m 0.287 0.381 0.384 6.905 7.202 7.196
pythia-70m 0.355 0.430 0.434 6.749 7.051 7.044
pythia-160m 0.454 0.522 0.526 6.877 6.969 6.961
pythia-410m 0.524 0.701 0.705 7.439 7.035 7.024
pythia-1.4b 0.574 0.799 0.803 8.254 7.325 7.312
pythia-2.8b 0.631 0.845 0.848 8.593 7.505 7.491
pythia-6.9b 0.637 0.855 0.859 8.538 7.373 7.360
pythia-12b 0.602 0.853 0.856 8.741 7.514 7.503
TinyLlama
TinyLlama-1.1B 0.641 0.762 0.767 7.802 7.081 7.065
Ours
ours-1.6b 0.710 0.778 0.782 8.563 7.712 7.694
ours-7b 0.784 0.837 0.839 8.982 8.028 8.016
ours-13b 0.661 0.807 0.810 8.584 7.674 7.665
(a) bloom (176b)
(b) gpt-neox-20b
(c) pythia-12b
(d) ours-13b
Figure 3: Relationship between the SMI metric and accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
Figure 4: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families.
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs (Vaswani et al., 2017). These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion (Scao et al., 2022).
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile (Black et al., 2021, 2022; Wang, 2021).
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable (Biderman et al., 2023b).
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions (Zhang et al., 2024; Soboleva et al., 2023; Li et al., 2023). The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
Evaluation setting
We evaluate these pre-trained models on the designated QA evaluation dataset. During inference, we configure the generation parameters with a temperature of 0.7, a maximum of 32 new tokens, and bfloat16 precision. A model response is considered correct if the expected object entity appears anywhere within the generated text. To obtain stable and reliable estimates of performance, we repeat the inference procedure 20 times for each of the 20 queries in the QA task, resulting in a total of 400 responses per model. The final accuracy is computed as the mean over all 400 responses. The full evaluation runs for approximately five days on 8 H100 GPUs.
The retrieval results from the pre-training corpus show that the co-occurrence frequency distribution of question and answer entities is highly skewed (see Figure 2). Most QA pairs co-occur between 10 and 1000 times, while only a small portion appears at very low or very high frequencies. To mitigate this imbalance and enable meaningful evaluation across different frequency ranges, we group QA data with similar metric values. For the co-occurrence and MI metrics, we compute average accuracy by binning values into intervals of 0.2 on the unnormalized scale. The SMI metric is then derived from the MI metric after binning.
Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship. Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
4.2 Experimental Results
Comparison of SMI, MI, and CO-OCCUR
Table 1 presents the R² and MSE results of CO-OCCUR, MI, and SMI across 24 models, with sizes ranging from 14M to 176B parameters. Overall, MI consistently improves over the co-occurrence baseline. For example, on gpt-neo-2.7B, R² increases from 0.672 (CO-OCCUR) to 0.829 (MI), while MSE decreases from 8.168 to 7.307. This suggests that jointly modeling frequency and specificity provides more informative signals for predicting closed-book QA performance. Incorporating model size into the metric (SMI) yields further gains across all models, though the magnitude of improvement is typically moderate. For instance, on pythia-12b, R² rises from 0.853 (MI) to 0.856 (SMI), and on bloom, from 0.764 to 0.766. These results suggest that incorporating model size into the metric provides a more principled account of the variance in observed QA accuracy.
Diminishing returns from data scaling
Despite the improved predictive power of SMI, we find that knowledge retention is limited by the nature of pre-training data. As shown in Figure 3, even when SMI reaches its maximum value of $1$, the closed-book QA accuracy does not exceed moderate levels. Specifically, the empirical upper bounds of accuracy for bloom, gpt-neox-20b, pythia-12b, and ours-13b are approximately $0.527$, $0.578$, $0.588$, and $0.630$, respectively. These results indicate that simply scaling the data cannot indefinitely improve factual recall.
Diminishing returns from model scaling
We further analyze the impact of model size on knowledge retention. Figure 4 shows that the average accuracy on the full evaluation set scales approximately linearly with the logarithm of model parameters, which is consistent with prior scaling-law observations. However, the growth rate becomes increasingly shallow for very large models: even at 176176B parameters, the accuracy remains below $0.2$. This finding suggests that model scaling alone is insufficient for achieving high factual coverage and further motivates the exploration of complementary approaches, such as retrieval-augmented generation or external memory modules.
5 Conclusion
In this work, we introduce a framework for quantifying knowledge retention in pre-trained language models, defined as the probability that a model reproduces factual information from its training corpus in a closed-book setting. We propose SMI, an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size, and demonstrate its effectiveness in forecasting closed-book QA accuracy. Our large-scale evaluation across 24 models shows that SMI achieves high predictive correlation (R² > 0.7 for models above 1B parameters) and remains robust across model families and domains.
Beyond prediction, our analysis reveals diminishing returns from scaling data and model size, suggesting an upper bound on factual knowledge that can be stored through pre-training alone. At the same time, SMI provides a diagnostic lens for data and model design: increasing the frequency of subject–object co-occurrences can strengthen knowledge memorization, filtering out irrelevant material can improve specificity, and scaling model size can further enhance retention, albeit with diminishing returns. These insights suggest that complementary strategies such as retrieval-augmented generation, external memory modules, and targeted fine-tuning may be necessary to further improve factual recall in LLMs.
We hope that our work provides a foundation for more systematic investigations into the interaction between corpus composition, model size, and factual knowledge retention, and informs future efforts to build models that combine parametric and non-parametric memory in a principled manner.
6 Limitations
While our study provides an information-theoretic framework for predicting knowledge retention, several limitations remain. First, our analysis is restricted to closed-book QA tasks and does not cover open-ended generation, reasoning-intensive tasks, or multi-hop factual queries, which may involve different memorization dynamics. Second, we rely on publicly available corpora and approximate document matching, which may introduce noise when measuring fact frequency and specificity. Third, our experiments focus on models with up to tens of billions of parameters, leaving open the question of whether the observed trends hold for frontier-scale models with hundreds of billions of parameters. Finally, our approach does not address how training dynamics, such as optimization schedules or curriculum learning, may interact with knowledge retention, which we leave for future work.
Appendix A AI Assistants in Research or Writing
In preparing this manuscript, AI assistants were employed solely to assist with refining the clarity, style, and readability of certain text segments. They were not involved in designing the study, developing or implementing the methodology, collecting or analyzing data, or generating the primary scientific contributions. All substantive research decisions, analyses, and conclusions are fully the responsibility of the authors.
Appendix B Licenses and Terms of Use
We use the following publicly available assets in this work:
We confirm that all licenses and terms of use are respected.
Appendix C Case Study
Table 2 highlights four counterintuitive cases in our pre-training data. From these, we observe the following: (1) In the first two cases, the subject and object rarely co-occur, suggesting a seemingly weak relationship. However, whenever the subject does appear, the object almost always appears alongside it, resulting in high accuracy. (2) In the last two cases, despite frequent co-occurrence between the subject and object, their individual occurrence frequencies are significantly higher, leading to low accuracy. For example, when prompted with “Paris is located in”, the model is more likely to respond with “France” rather than “Europe”.
These examples highlight the pivotal role of knowledge specificity in the memory of LLMs. It is not simply the frequencies of knowledge occurrence that enhance retention, but rather the specificity of the knowledge that proves more influential. This insight suggests that in downstream task training, even when domain-specific data is significantly fewer than the pre-training data, minimizing repetition with the pre-training data can increase the SMI, making it easier for LLMs to retain the relevant knowledge.
Appendix D Templates for Knowledge Triples
[…]
{ "plays": [ "[S] plays", "[S] performs", "[S] engages in playing", "[S] is an musician of", "[S] is involved in playing", "[S] is a performer of", "[S] executes", "[S] is a practitioner of", "[S] is an player of", "[S] is an artist of", "The genre that [S] performs is", "[S] captivates audiences by performing", "[S] performs the musical style of", "[S] is known for his performances in", "[S]’s musical expertise lies in", "[S] is recognized for playing", "[S] is a notable performer of", "[S]’s artistry is showcased in", "[S]’s musical inclinations are towards", "The musical category that [S] excels in is" ], "occupation": [ "The occupation of [S] is", "[S] works as an", "[S] is employed in the profession of an", "[S] pursues a profession as an", "[S] is a professional", "[S] is engaged in the profession of an", "[S]’s vocation is an", "[S] is an professional", "[S]’s role in life is that of an", "[S]’s professional identity is that of an", "The job of [S] is", "[S] is employed as an", "[S] has a career as an", "The profession of [S] is", "[S]’s line of work is", "[S] makes her career as an", "The career of [S] is", "[S] has chosen to be an", "[S] earns a living as an", "[S]’s professional life revolves around being an" ], "work": [ "[S] works in the field of", "[S] specializes in the field of", "[S]’s work focuses on", "[S] is engaged in the area of", "[S]’s professional activity lies in", "[S]’s expertise is in", "[S] operates within the field of", "[S]’s research is concentrated in", "[S] works extensively in", "[S] has dedicated his career to", "[S] is involved in the study of", "[S]’s field of study is", "[S] contributes to the discipline of", "[S] practices in the area of", "[S]’s work domain is", "[S] conducts his work in", "[S]’s field of expertise is", "[S]’s career revolves around", "[S] works professionally in", "[S]’s occupation relates to" ], "is-located": [ "[S] is located in", "The fatherland of [S] is in", "[S] can be found in", "[S] lives in", "[S] is situated within", "[S] is based in", "[S] resides in", "The country in which [S] lives is", "[S]’s location is in", "[S] stays in", "[S]’s address is in", "[S] is present in", "[S] spends most of his time in", "[S] belongs to", "[S]’s position is in", "You can find [S] in", "[S] is positioned in", "[S] exists in", "[S] stays primarily in", "The country in which [S] is located is" ], "locate": [ "[S] is located in", "The place in which [S] is held is", "[S] is held in", "The place where [S] takes place is", "[S] takes place in", "[S] is hosted in", "The venue for [S] is", "[S] happens in", "[S] is organized in", "[S] is staged in", "[S] occurs in", "The place for [S] is", "The location of [S] is", "[S] is based in", "[S] is set in", "[S]’s venue is in", "[S] is arranged in", "[S] is situated in", "People can take part in [S] in", "[S]’s setting is in" ], "found": [ "[S] was founded in", "[S] originated in", "[S] was established in", "[S] came into existence in", "[S] was formed in", "[S] had its beginnings in", "[S] started in", "[S] was created in", "[S] was initiated in", "[S] took shape in", "[S] was launched in", "[S] began in", "[S] emerged in", "The foundation of [S] took place in", "[S] originated from", "[S] was set up in", "[S] was first formed in", "[S] came to be in", "[S] had its origin in", "[S]’s formation occurred in" ], "born": [ "[S] was born in", "[S] came into the world in", "The birthplace of [S] is", "[S] entered life in", "[S] was delivered in", "[S]’s origin is", "[S] first saw the light in", "[S]’s birth occurred in", "[S] was brought into existence in", "[S]’s nativity is", "The place in which [S] was given birth to is", "[S] was born and raised in", "[S]’s roots are in", "The hometown of [S] is", "[S]’s cradle was in", "[S]’s entry into the world happened in", "[S]’s birth took place in", "[S] was born into the world in", "[S]’s life began in", "[S]’s starting point was" ], "die": [ "[S] died in", "[S] passed away in", "[S]’s death occurred in", "[S] lost his life in", "[S]’s demise happened in", "[S] perished in", "[S]’s end came in", "[S]’s life ended in", "[S] met his end in", "[S]’s passing was in", "[S]’s death took place in", "[S]’s final moments were in", "[S]’s last breath was in", "[S]’s life was lost in", "[S]’s expiration occurred in", "[S]’s fatality was in", "[S]’s death scene was in", "The place in which [S] passed away is", "[S]’s death location is", "[S]’s death site is" ], "air": [ "[S] was originally aired on", "[S] first aired on", "[S] premiered on", "[S] was initially broadcast on", "[S] debuted on", "[S] was first shown on", "[S] was originally broadcast on", "[S] was first transmitted on", "[S] was first telecast on", "[S] was first screened on", "[S] was first presented on", "[S] had its first airing on", "[S] was first televised on", "[S] was first released on", "[S] was first featured on", "[S] was first displayed on", "[S] was first exhibited on", "[S] was first put on air on", "[S] was first launched on", "[S] was first introduced on" ], "headquarter": [ "The headquarters of [S] is in", "[S]’s headquarters is in", "[S]’s main office is in", "[S]’s central office is in", "[S]’s head office is in", "The corporate office of [S] is in", "[S]’s principal office is in", "[S]’s administrative center is in", "[S]’s base of operations is in", "[S]’s HQ is in", "The headquarters’s location of [S] is in", "[S]’s headquarters office is in", "[S]’s primary office is in", "[S]’s main administrative office is in", "[S]’s central headquarters is in", "[S]’s main base is in", "The main hub of [S] is in", "[S]’s main building is in", "[S]’s headquarters is located in", "The place in which [S]’s headquarters is situated in" ], "capital": [ "[S] is the capital of", "[S] serves as the capital of", "[S] is recognized as the capital of", "[S] stands as the capital of", "[S] is known as the capital of", "[S] functions as the capital of", "[S] is officially the capital of", "[S] holds the title of capital of", "[S] is designated as the capital of", "[S] is the administrative capital of", "[S] is the political capital of", "[S] is the governmental capital of", "[S] is the principal capital of", "[S] is the main capital of", "[S] is the chief capital of", "[S] is the primary capital of", "[S] is the foremost capital of", "The country whose capital is [S] is", "[S] is known to the people as the capital of", "As the capital city, [S] has an extremely important place in" ], "citizen": [ "[S] is a citizen of", "The country where [S] holds citizenship and enjoys benefits is", "The nation that grants [S] the rights and privileges of citizenship is", "[S] is from the country of", "The place of [S]’s birth and upbringing is", "[S] belongs to", "[S]’s homeland is", "[S]’s home country is", "[S] is a resident of", "[S]’s origin is", "The land where [S]’s roots run deep is", "[S]’s place of origin is", "[S]’s country of origin is", "The nation that [S] calls home is", "[S]’s native country is", "[S]’s nationality is from", "The place where [S] was born and raised is", "[S]’s citizenship gives him the rights and privileges of a citizen in", "[S]’s citizenship makes him a part of the country’s community in", "The country that [S] resides in and holds citizenship is" ], "create": [ "[S] was created in", "The birthplace of [S] is", "[S] originated in", "The invention of [S] took place in", "[S] hails from", "The country that [S] was made in is", "[S] was initially developed in", "The country that introduced [S] is", "[S] was crafted in", "The place of origin for [S] is", "[S] was first produced in", "The homeland of [S] is", "The source of [S] is", "[S] was manufactured in", "The origin of [S]’s creation is", "[S] was designed in", "The place where [S] was conceived is", "[S] was made in", "The birth nation of [S] is", "[S] was developed in" ], "play": [ "[S] plays", "[S] is known for playing the", "[S]’s instrument is the", "[S] is a musician who plays the", "[S] is one who has mastered the art of the", "[S] has a gift for playing the", "Few can rival [S]’s prowess on the", "[S] is a true virtuoso of the", "[S]’s music is infused with the soulful sound of the", "Audiences are captivated by [S]’s", "[S]’s musical talents shine through his", "[S] is a skilled", "One can hear [S]’s passion in every", "With every note, [S] demonstrates his mastery of the", "[S] is a", "[S] plays the", "[S] specializes in the", "[S] performs on the", "Music flows through [S] and his", "As for instruments, [S] prefers the" ], "is-locate": [ "[S] is located in", "[S] lies in the", "[S] lies within the boundaries of", "The place where you can find [S] is", "[S] is situated in", "[S] is positioned in", "[S] exists in", "[S] is placed in", "[S] resides in", "The place where [S] can be fully appreciated is", "[S] can be found in", "[S] is in the area of", "[S] is part of the landscape of", "[S] is in the lands of", "The region that holds [S] is", "The location of [S] is", "[S] is in the sector of", "The place that [S] is a part of is", "[S] is in the section of", "[S] is in the region of" ] }